# Rheological Analysis and Evaluation of Measurement Techniques for the Curing Polymethylmethacrylate Bone Cement in Vertebroplasty

### v1

@@ Abstract
Vertebroplasty is a minimally invasive surgical procedure used to treat vertebral fractures, which conventionally involves injection of a polymethylmethacrylate (PMMA) bone cement into the fractured vertebra. A common risk associated with vertebroplasty is that of cement leaking out of the vertebra during the injection. This could happen since the flow behaviour of the bone cement is not conventional, but it varies with many factors like curing time, shear rate, and temperature. Therefore, experiments to quantify the cement’s flow properties are necessary to be able to understand and properly handle the bone cement. In this study, we aimed to characterize the behaviour of the PMMA bone cement in its curing stages, in order to obtain parameters that govern the flow behaviour during injection flow. The measurements were done using rotational and oscillatory rheometry as well as using a custom-made injetor setup replicating a typical vertebroplasty setting. Our results showed that the complex viscoelastic behaviour of the bone cement, which evolves over a two-phase curing process, can be significantly affected by deformations and temperature. We found that the results from rotational tests were affected by wall slip and “ridge”-like formations in the test sample, which we confirmed visually by microscopical assessment. We also found that the Cox-Merz rule to be conditionally valid, which impacts the use of oscillatory tests to obtain shear-thinning characteristics for the bone cement.
@@ 1 Introduction
Vertebral fractures are a major source of suffering and disability, carrying risk of back pain, immobility, and potentially severe post-complications that impact the quality of life 1, 2. Up to 28% of patients with osteoporotic fracture-related issues resulted in death by 2008 3. Unlike many other musculoskeletal complications where bone grafting or alternative methods can be used 4, the most commonly used procedure to treat vertebral fractures is vertebroplasty 5, 6, 7. In this procedure, a rapidly curing bone cement is injected into the porous cancellous part of the vertebra to serve as mechanical support 8, 9, 10 and allow early mobilization, especially in the elderly, where lack of mobility often causes a cascade of medical issues and often death. A successful vertebroplasty can provide instant pain relief 11 to the patient. On the other hand, improper injection could lead to cement leaking outside the vertebra, which could cause serious complications like pulmonary embolism and paralysis 12, 13. The bone cement is a material with complex mechanical behaviour dependent on many factors, which makes it hard to predict its behaviour during injection, especially for practitioners who rely on intuition, skill, and sensory feedback. Therefore, characterization of the curing bone cement using various measurement techniques is helpful to understand its behaviour during injection in various conditions. Mathematical constitutive models derived from the characterization data can quantify the various dependencies, which can be used in analytical or computational models to reliably predict the behaviour of the bone cement during injection 14, 15, 16, 17, 18, 19.
@@
The bone cements typically consist of ceramic components and fast-curing, minimally-toxic polymers that rapidly polymerize into, e.g. poly(methyl methacrylate) (PMMA) 8, 20, 21, 22. The rapid polymerization (curing) of bone cement is designed to provide the required structural support for the damaged vertebra, whereas ceramic components like zirconium dioxide function as radio-opacifiers that allow for easy visualization during the vertebroplasty procedure 23. Hydroxyapatite, a calcium phosphate mineral naturally found in bone, is also sometimes added to improve biocompatibility and to promote bone growth around the cement. The curing typically relies on free radical polymerization, the initiators react to form free radicals which cause the methlyl methacrylate (MMA) monomer molecules to form PMMA polymer chains 24, 25. The polymerization is an exothermic process 24, meaning that it is further accelerated by the heat released and the energy introduced, including mechanical shear.
@@
In general, it should be noted that PMMA bone cement is far from behaving like an ideal fluid, rather it has a dough-like consistency while it is curing. The behaviour of the bone cement can be categorized as viscoelastic, i.e. it is a combination of solid-like and fluid-like behaviour. Moreover, the behaviour is dependent on many factors, including time, flow conditions (e.g. shear rate), and temperature. Due to this multifactorial dependency, the complete characterization of the bone cement is a challenge. There are many previous studies measuring the rheological properties of the bone cements. Different studies use different brands of PMMA bone cement for their measurements. Moreover, the set of conditions used to carry out the measurements are different in each study. Not only the conditions, but also the methods used for the measurements are different in each study, e.g. while some studies report steady shear viscosity from rotational shear measurements, others report complex viscosity from oscillatory measurements. The steady state and complex viscosities can be compared only using the Cox-Merz rule 26, which says that the complex viscosity measured must be equal to the steady shear viscosity if the angular frequency is equal to the shear rate, respectively. However, this rule is not valid for all kinds of materials. Nevertheless, its validity for the PMMA bone cement is often assumed rather than being explicitly investigated. The differences in some of these studies are shown in Table 1. Due to these reasons, it is difficult to compare or compile the data from these studies to understand the complete rheological behaviour of a bone cement over the range of shear rates and temperatures expected during a vertebroplasty.
@@
Hence, the aim of this work was to understand the flow behaviour of the bone cement during injection in its curing phase in conditions relevant to vertebroplasty, taking Vertecem V+ as an example. The main objectives of this work were as follows:
@@
To quantify various aspects affecting injection flow behaviour, viz. viscosity, yield stress, loss factor, flow point, etc. from measurements
@@
To fit the experimental results to mathematical models that describe the rheological behaviour, like the power law model, and obtain the model parameters that can be used for prediction and simulation of injection flow behaviour
@@
To compare the different methodologies used for rheological measurement, viz. injection tests as well as rotational and oscillatory measurements on the rheometer, namely their advantages and limitations
@@
| Authors | Cement | Methods | Measurements and conditions |
| Boger et al. 27 | Unnamed | Oscillatory | Time evolution, 1 Hz, various |
| Deusser et al. 28 | Vertecem | Oscillatory | Time evolution, 1 Hz, 23 °C and 37 °C |
| Dunne and Orr 29 | Unnamed | Capillary | Time evolution |
| Farrar and Rose 30 | Various | Oscillatory | Time evolution, 0.05%, 5 Hz, various |
| Kolmeder and Lion 31 | Osteopal | Rotational CP | Time evolution, various |
| Capillary | Time evolution, various | ||
| Kolmeder et al. 32 | Unnamed | Oscillatory PP | Time evolution, 1 Hz, 0.1%, various |
| Krause et al. 33 | Various | Rotational CP | Time evolution, 0.04 and 1 s |
| Capillary | Shear rate dependence, 23 °C | ||
| Lepoutre et al. 17 111These studies assume the validity of the Cox-Merz rule. | Osteopal | Oscillatory PP | Time evolution, 50%, various and |
| Injector | Shear rate dependence, various | ||
| Lewis and Carroll 34 | Orthoset | Oscillatory CuP | Time evolution, 1 Hz, 18 °C |
| Lian et al. 18 111These studies assume the validity of the Cox-Merz rule. | Simplex | Oscillatory PP | Frequency sweep, 0.2% |
| Nicholas et al. 35 | Various | Oscillatory PP | Time evolution, 5 Hz |
@@ 2.1 Bone cement
The bone cement used for this study was purchased as non-sterile bulk material from OSARTIS GmbH (Germany), which is equivalent of the bone cement commercially sold by the name VertecemTM V+ by DePuy Synthes. The bone cement is supplied as a powder consisting of 40% zirconium dioxide, 15% hydroxyapatite, 44.6% poly(methyl methacrylate) (PMMA), and 0.4% benzoyl peroxide; and a liquid consisting of 99.35% methyl methacrylate (MMA) stabilized with 60 ppm hydroquinone, and 0.65% N,N-dimethyl-p-toluidine (DMPT). Typical application of the bone cement uses the provided cement mixing kit (DePuy Synthes), where the initiator and powder are mixed with MMA monomers in 26 grams of powder to 10 mL of liquid ratio, upon which immediately the user is required to push and pull the mixing chamber kit handle from endpoint to endpoint for 20 seconds, at a steady pace of 1–2 strokes per second. The resultant mixture has a dough-like consistency which can be filled into syringes and injected. The polymerization occurs through the free radical polymerization method, wherein the activator DMPT reacts with the initiator benzoyl peroxide to form free radicals. These free radicals react with monomer MMA molecules that cause them to form polymer chains of PMMA. PMMA is the main component of bone cement and is responsible for its mechanical properties. The entire polymerization process takes about 20 minutes at room temperature.
@@ 2.2 Rheometer measurements
The rheological measurements were performed on MCR302 rheometer from Anton Paar using a parallel plate (PP) setup. Disposable top and bottom plates were used for the measurements. The diameter of the top plate was 25 mm. The bone cement contains particles of size 5.2 ± 1.5 microns 36, hence the gap between the two plates was kept at 1.5 mm to ensure sufficiently larger gap size compared to the particle size 30 and avoid any particle-scale effects. The temperature during the measurements was controlled using the Peltier module of the rheometer, so that the measurements are not affected by the heat released during curing. The temperature was fixed at 23 °C in all tests, unless otherwise specified. The rheometer and the associated instrumentation was switched on for at least one hour to equilibrate the temperature and all other conditions before carrying out measurements.
The bone cement prepared using the typically provided mixing kit would produce a large offset of unnecessary material for measurements. Hence, we established a mixing process for our laboratory conditions that produced only sufficient amount of bone cement for each measurement. To ensure that our laboratory mixing process replicates the bone cement behaviour from the original mixing process, we established a benchmark test as quality control. In the test, the bone cement sample was subjected to oscillations at a constant torque of 3 mNm and 1 Hz frequency at 23 °C, with a measuring point every 5 seconds. After multiple trials, the successful and reproducible method consisted of the following steps:
@@
After 20 s, part of sample was gently dropped onto rheometer bottom plate and top plate was gently lowered down to avoid any disturbance.
@@
The test on the rheometer commenced once the bone cement was prepared and placed between the parallel plates. The time since start of mixing to start of test for all tests varied in the range of 180 30 seconds. For the characterization of the bone cement, various tests were done using the rheometer, namely
@@
Test Rh1a: Oscillatory test at 0.2% strain amplitude and 1 Hz frequency for 30 minutes: this test was done to observe the evolution of the bone cement properties in time at low strain value.
@@
Test Rh1b: Oscillatory test at 20% strain amplitude and 1 Hz frequency for 30 minutes: this test was done to observe the evolution of the bone cement properties in time under continuous larger deformation.
@@
Test Rh2a: Oscillatory test with 0.1% strain amplitude, 0.1 Hz (5 mins) – 20% strain amplitude, 1 Hz (3 mins) – 0.1% strain amplitude, 0.1 Hz (22 mins). This procedure was done to replicate a typical injection, in which the bone cement is at rest at first after mixing, then it is deformed as it is filled and injected through the syringe and cannula, and finally is at rest once it is injected. The aim here was to investigate how the bone cement breaks under deformation and recovers thereafter.
@@
Test Rh2b: Oscillatory test with 0.1% strain amplitude, 0.1 Hz, 23 °C (5 mins) – 20% strain amplitude, 1 Hz, 23 °C (3 mins) – 0.1% strain amplitude, 0.1 Hz, 37 °C (22 mins). This procedure was done to replicate a typical injection into a human vertebra, hence the change in temperature in the last step from room temperature to physiological body temperature.
Test Rh3: Rotational shear stress applied in steps of 0 Pa (1 min) - 100 Pa (1 min) - 0 Pa (1 min) - 500 Pa (1 min) - 0 Pa (1 min) - 2000 Pa (1 min) - 0 Pa (1 min) to evaluate creep behaviour.
Test Rh4: Rotational shear strain applied in steps of 0% (2 min) - 1% (2 min) - 0% (2 min) and 0% (2 min) - 100% (2 min) - 0% (2 min) to evaluate stress relaxation behaviour.
@@
Test Rh5: Amplitude sweep from 0.01% to 1000% strain amplitude, 1 Hz frequency to obtain the yield stress and the crossover point of storage and loss moduli.
@@
Test Rh6: Frequency sweeps from 0.1 Hz to 100 Hz at various strain amplitudes. The frequency sweep yields the dependence of the complex shear modulus and the complex viscosity on oscillation frequency. The strain amplitudes were chosen a posteriori as per the different regions obtained from the amplitude sweep test.
@@
Test Rh7a: Rotational shear rate sweep, various tests were done to obtain flow curves with the shear rate ranging over 2 or 3 orders of magnitude in each test. In this way, steady-state viscosity values for shear rates from 10 s to 10 s were obtained.
@@
Test Rh7b: Rotational shear rate sweep from 0.1 to 100 s with 100-grit sandpaper stuck to the top and bottom plates using double-sided tape. The test was done to check if there was a difference in the result with and without sandpaper, which would provide evidence for the occurrence of wall slip during the test.
@@
Test Rh8: Rotational tests with constant shear rate for 20 minutes. These tests are done at shear rates 10, 10, 10, 10, 1, and 100 s. These tests were done to evaluate the time-evolution of viscosity at various flow rates.

### v2

@@ Abstract
Vertebroplasty is a minimally invasive surgical procedure used to treat vertebral fractures, which conventionally involves injecting polymethylmethacrylate (PMMA) bone cement into the fractured vertebra. A common risk associated with vertebroplasty is cement leaking out of the vertebra during the injection, which may occur due to a lack of understanding of the bone cement’s complex flow behavior. Therefore, experiments to quantify the cement’s flow properties are necessary for understanding and proper handling of the bone cement. In this study, we aimed to characterize the behavior of the PMMA bone cement in its curing stages to obtain parameters that govern the flow behavior during the injection. We used rotational and oscillatory rheometry for our measurements, as well as a custom-made injector setup that replicated a typical vertebroplasty setting. Our results showed that the complex viscoelastic behavior of the bone cement is significantly affected by deformations and temperature. We found that the results from rotational tests, often used for characterizing the bone cement, are susceptible to measurement artifacts caused by wall slip and “ridge”-like formations in the test sample. We also found the Cox-Merz rule to be conditionally valid, which affects the use of oscillatory tests to obtain shear-thinning characteristics for the bone cement. Our findings identify important differences in the measured flow behavior of PMMA bone cement when assessed by different rheological methods, understanding which is crucial for its risk-free usage in downstream medical applications.
@@ 1 Introduction
Vertebral fractures are a major source of suffering and disability, carrying risk of back pain, immobility, and potentially severe post-complications that impact the quality of life 1, 2. Up to 28% of patients with osteoporotic fracture-related issues resulted in death by 2008 3. Unlike many other musculoskeletal complications where bone grafting or alternative methods can be used 4, the most commonly used procedure to treat vertebral fractures is vertebroplasty 5, 6, 7. In this procedure, a rapidly curing bone cement is injected into the porous cancellous part of the vertebra to serve as mechanical support 8, 9, 10 and allow early mobilization, especially in the elderly, where lack of mobility often causes a cascade of medical issues and often death 11. A successful vertebroplasty can provide instant pain relief 12 to the patient. On the other hand, improper injection could lead to cement leaking outside the vertebra, which could cause serious complications like pulmonary embolism and paralysis 13, 14. Bone cement is a material with complex rheological behavior dependent on many factors, which makes it hard to predict its behavior during injection, especially for practitioners who rely on intuition, skill, and sensory feedback. Therefore, characterization of the curing bone cement using various measurement techniques is helpful in understanding its behavior during injection in various conditions. Mathematical constitutive models derived from the characterization data can quantify the various dependencies, which can be used in analytical or computational models to reliably predict the behavior of the bone cement during injection 15, 16, 17, 18, 19, 20.
@@
The bone cement typically consists of ceramic components and fast-curing, minimally-toxic polymers that rapidly polymerize into, e.g. poly(methyl methacrylate) (PMMA) 8, 21, 22, 23. The rapid polymerization (curing) of bone cement is designed to provide the required structural support for the damaged vertebra, whereas ceramic components like zirconium dioxide function as radio-opacifiers that allow for easy visualization during the vertebroplasty procedure 24. Hydroxyapatite, a calcium phosphate mineral naturally found in bone, is also sometimes added to improve biocompatibility and promote bone growth around the cement. The curing typically relies on free radical polymerization, i.e. the initiators react to form free radicals, which cause the methyl methacrylate (MMA) monomer molecules to form PMMA polymer chains 25, 26. The PMMA polymerization is an exothermic process 25, that is known to increase local temperatures and contribute to some surrounding tissue damage. At the same time, increased local temperature increases the polymerization rate of PMMA, thereby generating a self-induced acceleration of the whole process 27.
| Authors | Cement | Methods | Characterization type |
| and conditions | |||
| Boger et al. 28 | Vertebroplastic | Oscillatory | Time, 1 Hz, various |
| Deusser et al. 29 | Vertecem | Oscillatory | Time, 1 Hz, 23 °C and 37 °C |
| Dunne and Orr 30 | Unnamed | Capillary | Time |
| Farrar and Rose 31 | Various | Oscillatory | Time, 0.05%, 5 Hz, various |
| Kolmeder and Lion 32 | Osteopal | Rotational CP | Time, various |
| Capillary | Time, various | ||
| Kolmeder et al. 33 | Unnamed | Oscillatory PP | Time, 1 Hz, 0.1%, various |
| Krause et al. 34 | Various | Rotational CP | Time, 0.04 and 1 s-1 |
| Capillary | Shear rate, 23 °C | ||
| Lepoutre et al. 18 * | Osteopal | Oscillatory PP | Time, 50%, various and |
| Injector | Shear rate, various | ||
| Lewis and Carroll 35 | Orthoset | Oscillatory CuP | Time, 1 Hz, 18 °C |
| Lian et al. 19 * | Simplex | Oscillatory PP | Frequency sweep, 0.2% |
| Nicholas et al. 36 | Various | Oscillatory PP | Time, 5 Hz |
@@
In general, it should be noted that PMMA bone cement is far from behaving like an ideal fluid, rather it has a dough-like consistency while it is curing. The behavior of the bone cement can be categorized as viscoelastic, i.e. it is a combination of solid-like and fluid-like behavior. Moreover, the behavior is dependent on many factors, including time, flow conditions (e.g. shear rate), and temperature. Due to this multifactorial dependency, the complete characterization of the bone cement is a challenge. Many previous studies have measured the rheological properties of the bone cements. Different studies use different brands of PMMA bone cement for their measurements. Moreover, the set of conditions used to carry out the measurements is different in each study. Not only the conditions but also the methods used for the measurements are different in each study, e.g. while some studies report steady shear viscosity from rotational shear measurements, others report complex viscosity from oscillatory measurements. The differences in some of these studies are shown in Table 1. Due to these reasons, it is difficult to compare or compile the data from these studies to understand the complete rheological behavior of a bone cement over the range of shear rates and temperatures expected during vertebroplasty. The steady-state and complex viscosities can be compared only using the Cox-Merz rule 37. The Cox-Merz rule is an empirical relationship indicating that for unfilled polymers the shear viscosity dependency on the shear rate can be predicted from the angular frequency dependency of the complex viscosity. There are studies in the literature that show that the Cox-Merz rule is applicable for PMMA 38. However, the many additives in the bone cement make it a highly filled polymer, and similar materials have been reported to not obey the Cox-Merz rule owing to their high molecular weight and complex intermolecular binding phenomena 39. Nevertheless, its validity for the PMMA bone cement is assumed in many studies rather than being explicitly investigated.
@@
Hence, the aim of this work was to understand the flow behavior of the bone cement during injection in its curing phase in conditions relevant to vertebroplasty, taking Vertecem V+ as an example. The main objectives of this work were as follows:
@@
To quantify various aspects affecting injection flow behavior, namely viscosity, yield stress, loss factor, flow point, etc. from measurements
@@
To fit the experimental results to mathematical models that describe the rheological behavior, like the power law model, and obtain the model parameters that can be used for prediction and simulation of injection flow behavior
@@
To compare the different methodologies used for rheological measurement, namely injection tests as well as rotational and oscillatory measurements on the rheometer, namely their advantages and limitations
@@ 2.1 Bone cement
The bone cement used for this study was purchased as non-sterile bulk material from OSARTIS GmbH (Germany), which is equivalent to the bone cement commercially sold by the name VertecemTM V+ by DePuy Synthes. The bone cement is supplied as a powder consisting of 40% zirconium dioxide, 15% hydroxyapatite, 44.6% poly(methyl methacrylate) (PMMA), and 0.4% benzoyl peroxide; and a liquid consisting of 99.35% methyl methacrylate (MMA) stabilized with 60 ppm hydroquinone, and 0.65% N,N-dimethyl-p-toluidine (DMPT), making it a highly filled polymer. Typical application of the bone cement uses the provided cement mixing kit (DePuy Synthes), where the initiator and powder are mixed with MMA monomers in 26 grams of powder to 10 mL of liquid ratio, upon which immediately the user is required to push and pull the mixing chamber kit handle from endpoint to endpoint for 20 seconds, at a steady pace of 1–2 strokes per second. The resultant mixture has a dough-like consistency which can be filled into syringes and injected. The polymerization occurs through the free radical polymerization method, wherein the activator DMPT reacts with the initiator benzoyl peroxide to form free radicals. These free radicals react with monomer MMA molecules causing them to form polymer chains of PMMA. PMMA is the main component of bone cement and is responsible for its mechanical properties. The entire polymerization process takes about 20 minutes at room temperature.
@@ 2.2 Rheometer measurements
The rheological measurements were performed on an MCR302 rheometer from Anton Paar using a parallel plate (PP) setup. Disposable top and bottom plates were used for the measurements. The diameter of the top plate was 25 mm. The bone cement contains particles of size 5.2 ± 1.5 microns 40, hence the gap between the two plates was kept at 1.5 mm to ensure a sufficiently larger gap size compared to the particle size 31 and avoid any particle-scale effects. The temperature during the measurements was controlled using the Peltier module of the rheometer, which maintained the temperature of the bottom plate. The temperature was fixed at 23 °C in all tests unless otherwise specified. The feedback constant temperature control of 23 °C ensures that the exothermic release of heat is not affecting our measured processes. The rheometer and the associated instrumentation were switched on for at least one hour to equilibrate the temperature and all other conditions before carrying out measurements.
The bone cement prepared using the typically provided mixing kit would produce a large offset of unnecessary material for measurements. On the other hand, the results of the rheological measurements are very sensitive to the mixing conditions. Hence, we established a mixing process for our laboratory conditions that not only produced just the necessary amount of bone cement for each measurement but also gave repeatable results on our benchmark test. In the test, the bone cement sample was subjected to oscillations at a maximum torque of 3 mNm and 1 Hz frequency at 23 °C, with a measuring point every 5 seconds. After multiple trials, the successful and reproducible method consisted of the following steps:
@@
After 20 seconds, part of the sample was gently dropped onto the rheometer bottom plate and the top plate was gently lowered down to avoid any disturbance.
@@
The results of the benchmark test are provided in the supporting information in Figure S1. The mixing method produced qualitatively reproducible results with quantitative deviations of up to 25%. Therefore, for the tests where capturing the bone cement response was sufficient, the measurements were not repeated to conserve material. For tests where the quantitative values were important, we did at least three repetitions of the tests, e.g. the amplitude and frequency sweep tests which were used to obtain the yield stress and the power law index respectively.
The bone cement used here does not have any waiting time 41, so it can be used or injected immediately after mixing. The test on the rheometer commenced once the bone cement was prepared and placed between the parallel plates. The time from the start of mixing to the start of test for all tests varied in the range of 182 33 seconds. For the characterization of the bone cement, various tests were done using the rheometer, the conditions of which were chosen in the context of the bone cement and vertebroplasty, based on information from previous studies in literature, like Krebs et al. 42, and analytical calculations. For the oscillatory tests, the complex modulus (norm of storage and loss modulus), the loss factor (the ratio of loss to storage modulus), and complex viscosity (norm of real and imaginary viscosity) were of main interest. For the rotational tests, the shear viscosity was the main quantity of interest. These tests are detailed below.
@@
Test Rh1a: In this test, the bone cement was subjected to oscillatory deformation of 0.2% strain amplitude and 1 Hz frequency for 30 minutes. The test was done to observe the evolution of the bone cement properties with time at a low strain value, i.e. nearly at rest condition.
@@
Test Rh1b: In this test, the bone cement was subjected to oscillatory deformation of 20% strain amplitude and 1 Hz frequency for 30 minutes. The test was done to observe the evolution of the bone cement properties in time under large deformations, i.e. in flow conditions.
@@
Test Rh2a: The test applied oscillatory deformation in three stages: (i) 0.1% strain amplitude, 0.1 Hz (5 mins) (ii) 20% strain amplitude, 1 Hz (3 mins) (iii) 0.1% strain amplitude, 0.1 Hz (22 mins). This procedure was done to replicate a typical injection, in which the bone cement is at rest at first after mixing, then it is deformed as it is filled and injected through the syringe and cannula, and finally is at rest after injection. The aim here was to investigate how the bone cement breaks under deformation and recovers thereafter.
@@
Test Rh2b: The test applied oscillatory deformation in three stages: (i) 0.1% strain amplitude, 0.1 Hz, 23 °C (5 mins) (ii) 20% strain amplitude, 1 Hz, 23 °C (3 mins) (iii) 0.1% strain amplitude, 0.1 Hz, 37 °C (22 mins). This procedure was the same as in test Rh2a but the temperature in the last stage was increased to human body temperature to replicate a typical injection into a human vertebra, and thereby investigate the recovery in those conditions.
Test Rh3: In this test, rotational shear stress was applied in steps of (i) 0 Pa (ii) 100 Pa (iii) 0 Pa (iv) 500 Pa (v) 0 Pa (vi) 2000 Pa (vii) 0 Pa, of one minute each, to evaluate creep behavior. The shear stress values were chosen a posteriori based on the yield stress obtained from the amplitude sweep test Rh5.
Test Rh4: Two tests were done, once where rotational shear strain was applied in steps of (i) 0% (ii) 1% (iii) 0%, of two minutes each, and another in steps of (i) 0% (ii) 100% (iii) 0%, of two minutes each, to evaluate stress relaxation behavior. The strain values were chosen a posteriori based on the yield stress obtained from the amplitude sweep test Rh5.
@@
Test Rh5: This was an amplitude sweep test, where the strain amplitude was gradually increased from 0.01% to 1000%. The test was done at 1 Hz frequency. The aim of the test was to obtain the yield stress, the crossover point, and the linear and non-linear viscoelastic ranges of the bone cement.
@@
Test Rh6: This was a frequency sweep test, where the frequency was gradually increased from 0.1 Hz to 100 Hz at various strain amplitudes. The strain amplitudes were chosen a posteriori as per the different regions obtained from the amplitude sweep test Rh5. The test aimed to obtain the dependence of the complex modulus and the complex viscosity on the oscillation frequency, and to use the complex viscosity-frequency curve for comparison to rotational shear rate sweep (test Rh7a) for checking the validity of the Cox-Merz rule.
@@
Test Rh7a: Various rotational shear rate sweep tests were done to obtain flow curves, i.e. viscosity change with shear rate, with each test spanning over two to three orders of magnitude of shear rate. In this way, steady-state viscosity values for shear rates from 10-4 s-1 to 103 s-1 were obtained to cover the possible range of shear rate during vertebroplasty.
@@
Test Rh7b: Rotational shear rate sweep from 0.1 to 100 s-1 were done with 100-grit sandpaper stuck to the top and bottom plates using double-sided tape. The test was done to check if there was a difference in the result with and without sandpaper, which would provide evidence for the occurrence of wall slip, i.e. insufficient friction between the rheometer plates and the bone cement sample for proper transfer of applied forces.
@@
Test Rh8: Various tests were done by applying deformation with constant rotational shear rate for 20 minutes. These tests were done at shear rates 10-4, 10-3, 10-2, 10-1, 1, and 100 s-1, to evaluate the time-evolution of viscosity at flow rates possible during vertebroplasty.